OpenStudy (anonymous):

One hundred students were asked whether they were taking psychology or biology. The responses showed that 61 were taking psychology, 18 were taking both, and 12 were taking neither. How many were taking biology?

OpenStudy (anonymous):

is the answer 45 taking biology?

OpenStudy (anonymous):

i got 27

OpenStudy (anonymous):

is that confirmed correct?

OpenStudy (anonymous):

i was the one asking the question cuz i didn't know for sure

OpenStudy (anonymous):

P=taking psychology; B=taking biology P or B = P + B - (P and B) 88 = 61 + B - 18 B=45

OpenStudy (anonymous):

there are 100 people to start

OpenStudy (anonymous):

61 are taking psycology and 12 aren't takin either so 73 aren't take biology

OpenStudy (anonymous):

27 are